# A new theory of continuum defects in terms of the general teleparallel geometry and exterior algebra

### v1

@@ 2.4 Path dependence of scalar product in a non-metric geometry
Again, the last two integrals are equal to zero in the second line because of the parallel transport rule of vectors. We used the first Bianchi identity (2a) $DQ_{ab}=R_{(ab)}$ at the last step. Consequently, because of existence of non-metricity parallel transport of the scalar product depends on the path. But, in teleparallel geometries defined by zero-curvature, i.e., both $R_{[ab]}=0$ and $R_{(ab)}=0$, it turns out to be independent of the path. This result is the same as the Proposition 2.1 of Ref.[7]. For more discussion on this issue one can consult [23] and the references therein.
@@ 3 Review of defect theory in the metric affine geometry
where $Q=Q^{a}{}_{a}$ and $\zeta^{c}=\zeta_{b}{}^{bc}$. Now, the four continuity conditions (37), (41), (43), (44) are simplified, respectively,
@@
For the metric compatible connection defined by the conditions $Q_{a}=0$ and $\zeta^{a}=0$, they turn out to be a much simpler forms
@@
These results are parallel to those of the Ref. [2],[3],[4]. Assuming $\alpha^{ab}$ and $\theta^{ab}$ to be small and of the same order, i.e. in the linear limit, we recover the well-known conservation laws or the kinematic equations of defects in a covariant formulation which are valid for even the non-Cartesian coordinates.
@@
In summary, we obtained the four equations of motion (37), (41), (43), (44) for defects (dislocation plus disclination plus metric anomaly) in the Riemann-Cartan-Weyl (or metric affine) geometry defined by the triple $(Q_{ab},T^{a},R^{a}{}_{b})$. Non-metricity 1-form is defined as density of metric anomalies, torsion 2-form is related with dislocation, anti-symmetric and symmetric components of full curvature 2-form with rotational disclination and metrical disclination, respectively. Physical picture of dislocation and rotational disclination are clear in material science, but visualization of metrical disclination and metric anomaly is blurry. Therefore, we intend to offer a new defect theory by narrowing the geometric realm.
4 A new theory of defects in the general teleparallel geometry
We remember the decomposition of the full connection, $\omega_{ab}=\widetilde{\omega}_{ab}+\mathrm{L}_{ab}$. Defect 1-form contains torsion (via contortion) and non-metricity, $\mathrm{L}_{ab}=K_{ab}+(\iota_{b}Q_{ac}-\iota_{a}Q_{bc})e^{c}+Q_{ab}$. Thus we realize that the anti-symmetric part is $\mathrm{L}_{[ab]}=K_{ab}+(\iota_{b}Q_{ac}-\iota_{a}Q_{bc})e^{c}:=\Omega_{ab}$ and the symmetric part is $\mathrm{L}_{(ab)}=Q_{ab}$.
@@ 4 A new theory of defects in the general teleparallel geometry
where $Q_{c}:=\iota_{c}Q=\delta^{ab}Q_{abc}$ and $C$ is a constant which is kept as a suitable unit conversion parameter. They satisfy the properties $\delta_{ab}\theta^{ab}=0$ and $P^{b}:=\iota_{c}Q^{bc}=\delta_{cd}{Q}^{bcd}=0$. The second equation can be rewritten as
@@
Thus, it is worthy to remark that when the non-metricity hence disclination vanishes, the general dislocation density reduces to the (pure) dislocation density $\alpha^{ab}=*(e^{a}\wedge T^{b})$ and $T^{a}=\alpha^{ba}*e_{b}$ as claimed by de Wit [11],[12],[13],[14]. The next step is to obtain continuity equations for the defect tensors. Thus, the crucial point is to compute the covariant exterior derivative of the equations (51a) and (54).
@@
These are our new continuity equations such as those (47).
@@ 5 Discussion
Since dislocation density and disclination density are useful quantities to obtain information about many elastic fields such as distortion, strain, dilatation, stress and rotation caused by defects in a continuum, it makes sense to look for a complete theory that explains their behavior. One of the various ways to do that is to adopt the geometric approach. As far as we know, in almost all studies conducted in this direction in the literature, the dislocation density is matched with the torsion tensor of the Riemann-Cartan geometry, and the disclination density with the full curvature tensor [1],[2],[3],[4],[5],[6],[7],[8]. For a summary of that formulation, one can look at the equations (47) and later of Section 3. Those arguments are based on the definitions given by equations (33), together with $Q_{ab}=0$, hence $R_{(ab)}=0$ and $\zeta_{abc}=0$ constraints. In that formulation, the geometric arena suitable for discussing the behaviour of only dislocation type defects appears naturally as metric (Weitzenböck) teleparallel geometry. Conversely, we are left with Riemannian geometry as the appropriate geometry for dislocation-free continuum.
Looking at that formulation from a different perspective will give us new insights. If one inserts the decomposition of full curvature given by the equation (9a) or (10), together with the metricity condition, into the definition (33b), it becomes clear that there are contributions to the disclination density from the torsion tensor via the contortion 1-form inside of $\mathrm{L}_{ab}$. Therefore, according to the pairing between torsion and dislocation density (33a), since they are proportional to each other, there are contributions coming from the dislocation density to the disclination density. However, this conclusion is the opposite of what de Wit says in the papers [11],[12],[13],[14]. As we stated in the introduction, we adopt de Wit’s terminology and approach regarding the continuum defects. According to de Wit’s articles, if there is dislocation in a continuum, it contains some of them in the (general) dislocation density. When disclination is eliminated, the general dislocation density turns into (pure) dislocation density. To achieve this phenomenological result, we matched the symmetric part of the defect 1-form with the disclination density, see (51). Accordingly, we also matched the dislocation density to the anti-symmetric part of the defect 1-form, see (53). These correspondences appear the first time in this work. Thus, a disclination-free continuum, $\theta_{ab}=0$ hence $Q_{ab}=0$, can be represented by again the metric (Weitzenböck) teleparallel geometry. The reverse case characterized by the configuration, $T^{a}=0$ hence $\alpha_{ab}=0$, but $\theta_{ab}\neq 0$, is not always possible because of the equation (53a). This result is in accordance with de Wit’s conclusion. The geometric realm of the latter case is the symmetric teleparallel geometry.
@@
The ground-state configurations of a continuum with defects can be determined by the extrema of a free-energy integral $F=\int_{B}L$ where $B$ is the body manifold corresponding to the three dimensional abstract manifold $M$ and $L$ is called the Lagrange 3-form [2],[3],[4],[5],[6]. One of our future projects is to write down some possible Lagrange 3-forms containing torsion and non-metricity, obtain variational field equations, find some classes of exact solutions and compare them with the results coming from the material science world. Thus, we hope that our new and strict mathematical model for material defects may predict new insights on the behaviour of material.
In this work we introduced a new theory of defect fields for static configurations. A natural extension of that is to consider its dynamical counterpart. By inclusion of the defect velocity tensor field similar to those in [24],[25] the results of our work could be generalized. This remains an open problem waiting to be solved.
@@ Appendix B Relation between disclination and non-metricity
where $A,B,C,K$ are arbitrary constants. The first relation satisfies the property $\delta_{ab}\theta^{ab}=0$. We denote $Q_{abc}=\iota_{c}Q_{ab}$ and $Q_{c}=\iota_{c}Q=\delta^{ab}Q_{abc}=Q^{a}{}_{ac}$ where $Q_{(ab)c}$ by definition. We also define the second trace $P_{a}=\delta^{bc}Q_{abc}=\iota^{b}Q_{ab}=Q_{ab}{}^{b}$ along with the first trace $Q_{c}=\delta^{ab}Q_{abc}$ of non-metricity. We compute the second trace from the equation (72b) as

### v2

@@ 2.3 Path independence of scalar product in the non-metric teleparallel geometries
Again, the last two integrals are equal to zero in the second line because of the parallel transport rule of vectors. We used the first Bianchi identity (2a) $DQ_{ab}=R_{(ab)}$ at the last step. Consequently, because of existence of non-metricity parallel transport of the scalar product seems depending on the path. But, in teleparallel geometries defined by zero-curvature, i.e., both $R_{[ab]}=0$ and $R_{(ab)}=0$, it turns out to be independent of the path. This result is the same as the Proposition 2.1 of Ref.[10]. For more discussion on this issue one can consult [33] and the references therein.
@@ 3 Review of defect theory in the metric affine geometry
where $\zeta^{c}=\zeta_{b}{}^{bc}$. Now, the four continuity conditions (27), (31), (33), (34) are simplified, respectively,
@@
For the metric compatible connection defined by the conditions $Q_{a}=0$ and $\zeta^{a}=0$, we lands to the Riemann-Cartan geometry and the continuity equations turn out to be a much simpler forms
@@
These results are parallel to those of the Ref. [3]-[5]. Assuming $\alpha^{ab}$ and $\theta^{ab}$ to be small and of the same order, i.e. in the linear limit, we recover the well-known conservation laws or the kinematic equations of defects in a covariant formulation which are valid for even the non-Cartesian coordinates.
@@
In summary, we obtained the four equations of motion (27), (31), (33), (34) for defects (dislocation plus disclination plus metric anomaly) in the Riemann-Cartan-Weyl (or metric affine) geometry defined by the triple $(Q_{ab},T^{a},R^{a}{}_{b})$. Non-metricity 1-form is defined as density of metric anomalies, torsion 2-form is related with dislocation, anti-symmetric and symmetric components of full curvature 2-form with rotational disclination and metrical disclination, respectively. Physical picture of dislocation and rotational disclination are clear in material science, but visualization of metrical disclination and metric anomaly is blurry. Therefore, we intend to offer a new formulation of defect theory by narrowing the geometric realm.
4 A new formulation of defect theory in terms of the general teleparallel geometry
It is worthy to state explicitly that this section is the original content of our paper. We remember the decomposition of the full connection, $\omega_{ab}=\widetilde{\omega}_{ab}+\mathrm{L}_{ab}$. Defect 1-form contains torsion (via contortion) and non-metricity, $\mathrm{L}_{ab}=K_{ab}+(\iota_{b}Q_{ac}-\iota_{a}Q_{bc})e^{c}+Q_{ab}$. Thus we realize that the anti-symmetric part is $\mathrm{L}_{[ab]}=K_{ab}+(\iota_{b}Q_{ac}-\iota_{a}Q_{bc})e^{c}:=\Omega_{ab}$ and the symmetric part is $\mathrm{L}_{(ab)}=Q_{ab}$.
@@ 4 A new formulation of defect theory in terms of the general teleparallel geometry
where $C$ is a constant which is kept as a suitable unit conversion parameter. They satisfy the properties $\delta_{ab}\theta^{ab}=0$ and $P^{b}=0$. The second equation can be rewritten as
@@
Thus, it is worthy to remark that when the non-metricity hence disclination vanishes, the general dislocation density reduces to the (pure) dislocation density $\alpha^{ab}=*(e^{a}\wedge T^{b})$ and $T^{a}=\alpha^{ba}*e_{b}$ as claimed by de Wit [21]-[24]. The next step is to obtain continuity equations for the defect tensors. Thus, the crucial point is to compute the covariant exterior derivative of the equations (41a) and (44).
@@
These are our continuity equations such as those (37). We want again to emphasize that although the Riemann-Cartan geometry in which standard geometric formulation of defect theory is established does not have a metric formulation, the general teleparallel geometry does have a metric formulation [28]. This result is one of motivations directing us to do this work.
@@ 5 Discussion
Since dislocation density and disclination density are useful quantities to obtain information about many elastic fields such as distortion, strain, dilatation, stress and rotation caused by defects in a continuum, it makes sense to look for a complete theory that explains their behavior. One of the various ways to do that is to adopt the geometric approach. As far as we know, in almost all studies conducted in this direction in the literature, the dislocation density is matched with the torsion tensor of the Riemann-Cartan geometry, and the disclination density with the full curvature tensor [2]-[11]. For a summary of that formulation, one can look at the equations (37) and later of Section 3. Those arguments are based on the definitions given by equations (23), together with $Q_{ab}=0$, hence $R_{(ab)}=0$ and $\zeta_{abc}=0$ constraints. In that formulation, the geometric arena suitable for discussing the behaviour of only dislocation type defects appears naturally as metric (Weitzenböck) teleparallel geometry. Conversely, we are left with Riemannian geometry as the appropriate geometry for dislocation-free continuum.
Looking at that formulation from a different perspective will give us new insights. If one inserts the decomposition of full curvature given by the equation (10a) or (11), together with the metricity condition, into the definition (23b), it becomes clear that there are contributions to the disclination density from the torsion tensor via the contortion 1-form inside of $\mathrm{L}_{ab}$. Therefore, according to the pairing between torsion and dislocation density (23a), since they are proportional to each other, there are contributions coming from the dislocation density to the disclination density. However, this conclusion is the opposite of what de Wit says in the papers [21]-[24]. As we stated in the introduction, we adopt de Wit’s terminology and approach regarding the continuum defects. According to de Wit’s articles, if there is dislocation in a continuum, it contains some of them in the (general) dislocation density. When disclination is eliminated, the general dislocation density turns into (pure) dislocation density. To achieve this phenomenological result, we matched the symmetric part of the defect 1-form with the disclination density, see (41). Accordingly, we also matched the dislocation density to the anti-symmetric part of the defect 1-form, see (43). These correspondences appear the first time in this paper. Thus, a disclination-free continuum, $\theta_{ab}=0$ hence $Q_{ab}=0$, can be represented by again the metric (Weitzenböck) teleparallel geometry. The reverse case characterized by the configuration, $T^{a}=0$ hence $\alpha_{ab}=0$, but $\theta_{ab}\neq 0$, is not always possible because of the equation (43a). This result is in accordance with de Wit’s conclusion. The geometric realm of the latter case is the symmetric teleparallel geometry.
@@
The ground-state configurations of a continuum with defects can be determined by the extrema of a free-energy integral $F=\int_{B}L$ where $B$ is the body manifold corresponding to the three dimensional abstract manifold $M$ and $L$ is called the Lagrange 3-form [3]-[9]. One of our future projects is to write down some possible Lagrange 3-forms containing torsion and non-metricity, obtain variational field equations, find some classes of exact solutions and compare them with the results coming from the material science world. Thus, we hope that our new and strict mathematical model for material defects may predict new insights on the behavior of material.
In this work we introduced a new formulation of defect fields for static configurations. A natural extension of that is to consider its dynamical counterpart. By inclusion of the defect velocity tensor field similar to those in [34],[35] the results of our work could be generalized. That generalization remains an open problem waiting to be solved.
@@ Appendix B Relation between disclination and non-metricity
where $A,B,C,K$ are arbitrary constants. The first relation satisfies the property $\delta_{ab}\theta^{ab}=0$. We denote $Q_{abc}=\iota_{c}Q_{ab}$ and $Q_{c}=\iota_{c}Q=\delta^{ab}Q_{abc}=Q^{a}{}_{ac}$ where $Q_{(ab)c}$ by definition. We also define the second trace $P_{a}=\delta^{bc}Q_{abc}=\iota^{b}Q_{ab}=Q_{ab}{}^{b}$ along with the first trace $Q_{c}=\delta^{ab}Q_{abc}$ of non-metricity. We compute the second kind trace from the equation (62b) as
@@
Acknowledgements: TD acknowledges partial support by the Turkish Academy of Sciences.
Data Availability Statement: No data associated in the manuscript.